# Toward a unified theory for common affine roots of general sets of multivariate polynomials

### v3

@@ 1 Introduction
For a univariate non-zero polynomial over an arbitrary field the number of roots is at most equal to its degree, and if the field is algebraically closed and if roots are counted with multiplicity then equality holds. This result is in close correspondence with the fact that the interpolation problem of finding a polynomial of lowest possible degree evaluating to zero on a finite point-set boils down to calculating the product of linear terms each having one of the desired elements as root, and by repeating terms one can also deal with multiplicity. The situation is very much in contrast to that of multivariate polynomials. Firstly, a multivariate polynomial ring is not a principal ideal domain, and therefore the considered interpolation problem can not in general be about multiplying linear factors. Secondly, a non-constant multivariate polynomial often has infinitely many roots and always does if the field is algebraically closed, and therefore if the aim is to consider a case of finitely many roots then one either needs to restrict to roots from a prescribed finite point-set or one must study mutual roots of a set of polynomials which together define a zero-dimensional ideal.
@@
Our contribution is to address the problem of common roots of general polynomials where the only assumption is that no two of them have the same leading monomial which is no restriction as we can always apply the division algorithm [10][Sec, 2, Par. 3]] without effecting the set of common roots. Similarly, we consider any finite point-sets without any restriction on their structure. Therefore, our answer to (II) should be compared to Proposition 1 and our answer to (I) should be compared to the Schwartz-Zippel bound, the Alon-Füredi bound and the generalized Alon-Füredi bound [7]. In both cases the material of the present paper can be viewed as improvements and of course also as enhancements as we in contrast to the mentioned bounds do not restrict to single polynomials but rather consider arbitrary finite sets of linearly independent polynomials.
In the present paper we give refined answers to both (I) and (II) taking the leading monomial into account and we enhance the methods to treat the case of common roots of a set of more polynomials. Knowing the leading monomial corresponds to having important information regarding the most crucial coefficient and thereby our insight addresses the remark by Tao. More importantly, we demonstrate a very strong relationship between our answers to (I) and (II) which is then our proposal for the requested useful unification, except we do not treat multiplicity.
@@ 2 Main results on multivariate polynomials
As a special case of Corollary 9 we shall obtain a natural counterpart to Theorem 5 in the case of finite Cartesian product point-sets $A_{1}\times\cdots\times A_{m}$. We shall later argue, that it is in a certain sense not a restriction to consider only such point-sets when formulating the counterpart to Theorem 5 (see Remark 11 below). For finite Cartesian product point-sets the $I$ in Corollary 9 becomes
@@
Let $A_{1}\times\cdots\times A_{m}$ be a finite point-set and consider an arbitrary monomial ordering $\prec$. Let $I$ and $\Delta_{\prec}(I)$, respectively, be as in (3) and (4), respectively. Consider pairwise different monomials $M_{1},\ldots,M_{t}\in\Delta_{\prec}(I)$. For general set of $k\leq t$ polynomials $F_{1},\ldots,F_{k}\in{\mathbb{F}}[X_{1},\ldots,X_{m}]$ with ${\mbox{Supp}}(F_{i})\subseteq\Delta_{\prec}(I)$, $i=1,\ldots,k$ and leading monomials being pairwise different all belonging to $\{M_{1},\ldots,M_{t}\}$ the number of non-common roots of $F_{1},\ldots,F_{k}$ in $A_{1}\times\cdots\times A_{m}$ is at least
@@
$=$ $\min\{\mu\left(\dfrac{N}{M_{i_{1}}},\ldots,\dfrac{N}{M_{i_{k}}}\right)\mid 1\leq i_{1}<\cdots<i_{k}\leq t\}$
where $N:=X_{1}^{\#A_{1}-1}\cdots X_{m}^{\#A_{m}-1}$. This bound is sharp.
@@
Clearly, for general $F_{1},\ldots,F_{k}$ as in Theorem 5 the smallest attainable number of non-common roots in $A_{1}\times\cdots\times A_{m}$ equals $M_{k}(C_{1},C_{2})$ which as demonstrated in the coming sections equals (5). In the coming sections we shall also demonstrate that if $A\subseteq A_{1}\times\cdots\times A_{m}$ satisfies $A<M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ then there exist $k$ linearly independent polynomials as in Theorem 5 which all evaluates to zero on $A$, but that for some $A\subseteq A_{1}\times\cdots\times A_{m}$ with $A=M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ this does not hold. Here, the dual space (or null-space) is with respect to the usual inner product. Finally, as we shall demonstrate in the next two sections of the paper the right hand side of (1) serves as a lower bound on $M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$. We conclude that problem (I) and problem (II) and our answers to them are basically identical, but applied to dual spaces. All claims of the present remark follows by inspection of the proof of Theorem 5 in Section 4.
@@
In [6] it is already noted that the generalized Alon-Füredi bound is a consequence of the footprint bound. We here give some more details including discussing when the two bounds may or may not coincide. Consider a graded monomial ordering, i.e. a monomial ordering $\prec$ satisfying that if $\deg N_{1}<\deg N_{2}$ then necessarily $N_{1}\prec N_{2}$. There are many such orderings, including $m!$ degree lexicographic ordering where in addition to the above property ties are broken by the rule that if $K_{1}$ and $K_{2}$ are of the same degree then $K_{1}\prec K_{2}$ if $K_{1}\prec_{lex}K_{2}$ for the given chosen lexicographic ordering. For a graded monomial ordering for sure the leading monomial ${\mbox{lm}}(F)$ belongs to the set considered in (7) and Theorem 12 follows imediately from Corollary 10.
@@ 3 Results in linear algebra setting
We next introduce the concept of one-way well-behaving pairs which is more general than the well-known concepts of well-behaving pairs and weakly well-behaving pairs [36] allowing for a more powerful description [28]. Consider an additional ordered basis $B^{\prime}=\{\vec{b}^{\prime}_{1},\ldots\vec{b}^{\prime}_{n}\}$ (which may or may not be equal to $B$). Let ${\mathcal{I}}:=\{1,2,\ldots,n\}$. An ordered pair $(i,j)\in{\mathcal{I}}^{2}$ is said to be one-way well-behaving (OWB) if $\bar{\rho}(\vec{b}_{u}\ast\vec{b}^{\prime}_{j})<\bar{\rho}(\vec{b}_{i}\ast\vec{b}^{\prime}_{j})$ for $u<i$. Given bases $B,B^{\prime}$ as above consider for $l=1,2,\ldots,n$ and $i=1,2,\ldots,n$ the following sets
@@ 4 Proof of Theorem 5
$C_{2}^{\perp}$ $=$ ${\mbox{ev}}\left(({\mathcal{L}}_{2}^{\perp}:={\mbox{Span}}_{\mathbb{F}_{q}}\{M\mid M\in\Box,M\preceq\dfrac{N}{M_{1}}\})+I\right)$
$C_{1}^{\perp}$ $=$ ${\mbox{ev}}\left(({\mathcal{L}}_{1}^{\perp}:={\mbox{Span}}_{\mathbb{F}_{q}}\{M\mid M\in\Box,M\prec\dfrac{N}{M_{t}}\})+I\right)$
where $N:=X_{1}^{q-1}\cdots X_{m}^{q-1}$. Clearly ${\mbox{lm}}({\mathcal{L}}^{\perp}_{2}\backslash{\mathcal{L}}_{1}^{\perp})=\left\{\dfrac{N}{M_{t}},\ldots,\dfrac{N}{M_{1}}\right\}$ and the last part of Theorem 5 follows from the last part of Theorem 17 in combination with the last part of Corollary 10.
@@ 5 Concluding remarks
We further pose it as a research problem to enhance our findings to hold for integral domains following up on [7, 40] and possibly also to treat skew fields and other non-commutative algebraic structures where the OWB point of view in Section 3 may prove essential. Similarly it would be interesting to see if some of the methods in the present paper could be applied to treat projective points beyond what is already in the litterature e.g. in [12, 5, 37].
We note that like other recent results on roots of sets of multivariate polynomials, e.g. [17], also the present paper is related to work in coding theory and cryptography. For this particular paper the inspiration came from work on ramp secret sharing schemes [24]. Elaborating further on this relation we observe, that Theorem 17 should have interesting implications for one-point algebraic geometric codes [43, Sec. II] and other code constructions where one can obtain information on the functions $\bar{\sigma}$ as well as $\bar{\mu}$.

### v4

@@ 1 Introduction
For a univariate non-zero polynomial over an arbitrary field the number of roots is at most equal to its degree, and if the field is algebraically closed and if roots are counted with multiplicity then equality holds. This result is in close correspondence with the fact that the interpolation problem of finding a polynomial of lowest possible degree evaluating to zero on a finite point-set boils down to calculating the product of linear terms each having one of the desired elements as root, and by repeating terms one can also deal with multiplicity. The situation is very much in contrast to that of multivariate polynomials. Firstly, a multivariate polynomial ring is not a principal ideal domain, and therefore the interpolation problem can not in general be about multiplying linear factors. Secondly, a non-constant multivariate polynomial often has infinitely many roots and always does if the field is algebraically closed, and therefore if the aim is to consider a case of finitely many roots then one either needs to restrict to roots from a prescribed finite point-set or one must study mutual roots of a set of polynomials which together define a zero-dimensional ideal.
@@
Our contribution is to address the problem of common roots of general polynomials where the only assumption is that no two of them have the same leading monomial which is no restriction as we can always apply the division algorithm [10][Sec, 2, Par. 3]] without effecting the set of common roots. Similarly, we consider any finite point-sets without any restriction on their structure. Therefore, our answer to (II) should be compared to Proposition 1 and our answer to (I) should be compared to the Schwartz-Zippel bound, the Alon-Füredi bound and the generalized Alon-Füredi bound [7]. In both cases the material of the present paper can be viewed as improvements and of course also as enhancements as we in contrast to the mentioned bounds do not restrict to single polynomials but rather consider arbitrary finite sets of linearly independent polynomials. We should mention that the problems (I) and (II) have also been thoroughly treated in [46], but their answers are of a completely different nature than ours. In contrast to them we on the one hand do not assume any structure on the point-set, but on the other hand rely on information regarding leading monomial rather than just degree.
In the present paper we give refined and enhanced answers to both (I) and (II) using the leading monomials of a set of polynomials as a measure on their sizes. Knowing the leading monomial corresponds to having important information regarding the most crucial coefficient and thereby our insight addresses the remark by Tao. More importantly, we demonstrate a very strong relationship between our answers to (I) and (II) which is then our proposal for the requested useful unification, except we do not treat multiplicity.
@@ 2 Main results on multivariate polynomials
As a special case of Corollary 9 we may consider the situation where the affine variety equals a finite Cartesian product point-set $A_{1}\times\cdots\times A_{m}$. As shall become clear in Remark 11 below a study of this case surprisingly holds the key to establish the connection between our answers to (I) and (II). For finite Cartesian product point-sets the $I$ in Corollary 9 becomes
@@
Let $A_{1}\times\cdots\times A_{m}$ be a finite point-set and consider an arbitrary monomial ordering $\prec$. Let $I$ be as in (3) which implies that $\Delta_{\prec}(I)$ equals (4). Consider pairwise different monomials $M_{1},\ldots,M_{t}\in\Delta_{\prec}(I)$. For general set of $k\leq t$ polynomials $F_{1},\ldots,F_{k}\in{\mathbb{F}}[X_{1},\ldots,X_{m}]$ with ${\mbox{Supp}}(F_{i})\subseteq\Delta_{\prec}(I)$, $i=1,\ldots,k$ and leading monomials being pairwise different all belonging to $\{M_{1},\ldots,M_{t}\}$ the number of non-common roots of $F_{1},\ldots,F_{k}$ in $A_{1}\times\cdots\times A_{m}$ is at least
@@
$=$ $\min\{\mu\left(\dfrac{K}{M_{i_{1}}},\ldots,\dfrac{K}{M_{i_{k}}}\right)\mid 1\leq i_{1}<\cdots<i_{k}\leq t\}$
where $K:=X_{1}^{\#A_{1}-1}\cdots X_{m}^{\#A_{m}-1}$. This bound is sharp.
@@
Clearly, for general $F_{1},\ldots,F_{k}$ as in Theorem 5 the smallest attainable number of non-common roots in $A_{1}\times\cdots\times A_{m}$ equals $M_{k}(C_{1},C_{2})$ which as demonstrated in the coming sections equals (5). In the coming sections we shall also demonstrate that if $A\subseteq A_{1}\times\cdots\times A_{m}$ satisfies $A<M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ then there exist $k$ linearly independent polynomials as in Theorem 5 which all evaluates to zero on $A$, but that for some $A\subseteq A_{1}\times\cdots\times A_{m}$ with $A=M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ this does not hold. Here, the dual space (or null-space) is with respect to the usual inner product. Finally, as we shall demonstrate in the next two sections of the paper the right hand side of (1) serves as a lower bound on $M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$. We conclude that problem (I) and problem (II) and our answers to them are basically identical, but applied to dual spaces. All claims of the present remark follow by inspection of the proof of Theorem 5 in Section 4.
@@
In [6] it is already noted that the generalized Alon-Füredi bound is a consequence of the footprint bound. We here give some more details including discussing when the two bounds may or may not coincide. Consider a graded monomial ordering, i.e. a monomial ordering $\prec$ satisfying that if $\deg N_{1}<\deg N_{2}$ then necessarily $N_{1}\prec N_{2}$. There are many such orderings, including $m!$ degree lexicographic orderings where in addition to the above property ties are broken by the rule that if $K_{1}$ and $K_{2}$ are of the same degree then $K_{1}\prec K_{2}$ if $K_{1}\prec_{lex}K_{2}$ for the given chosen lexicographic ordering. For a graded monomial ordering for sure the leading monomial ${\mbox{lm}}(F)$ belongs to the set considered in (7) and Theorem 12 follows immediately from Corollary 10.
@@ 3 Results in linear algebra setting
We next introduce the concept of one-way well-behaving pairs which is more general than the well-known concepts of well-behaving pairs and weakly well-behaving pairs [36] allowing for a more powerful description [28] with a wider range of possible future applications of our results. Consider an additional ordered basis $B^{\prime}=\{\vec{b}^{\prime}_{1},\ldots\vec{b}^{\prime}_{n}\}$ (which may or may not be equal to $B$). Let ${\mathcal{I}}:=\{1,2,\ldots,n\}$. An ordered pair $(i,j)\in{\mathcal{I}}^{2}$ is said to be one-way well-behaving (OWB) if $\bar{\rho}(\vec{b}_{u}\ast\vec{b}^{\prime}_{j})<\bar{\rho}(\vec{b}_{i}\ast\vec{b}^{\prime}_{j})$ for $u<i$. Given bases $B,B^{\prime}$ as above consider for $l=1,2,\ldots,n$ and $i=1,2,\ldots,n$ the following sets
@@ 4 Proof of Theorem 5
$C_{2}^{\perp}$ $=$ ${\mbox{ev}}\left(({\mathcal{L}}_{2}^{\perp}:={\mbox{Span}}_{\mathbb{F}_{q}}\{M\mid M\in\Box,M\preceq\dfrac{K}{M_{1}}\})+I\right)$
$C_{1}^{\perp}$ $=$ ${\mbox{ev}}\left(({\mathcal{L}}_{1}^{\perp}:={\mbox{Span}}_{\mathbb{F}_{q}}\{M\mid M\in\Box,M\prec\dfrac{K}{M_{t}}\})+I\right)$
where $K:=X_{1}^{q-1}\cdots X_{m}^{q-1}$. Clearly ${\mbox{lm}}({\mathcal{L}}^{\perp}_{2}\backslash{\mathcal{L}}_{1}^{\perp})=\left\{\dfrac{K}{M_{t}},\ldots,\dfrac{K}{M_{1}}\right\}$ and the last part of Theorem 5 follows from the last part of Theorem 17 in combination with the last part of Corollary 10.
@@ 5 Concluding remarks
We further pose it as a research problem to enhance our findings to hold for integral domains following up on [7, 40] and possibly also to treat skew fields and other non-commutative algebraic structures where the OWB point of view in Section 3 may prove essential. Similarly it would be interesting to see if some of the methods in the present paper could be applied to treat projective points beyond what is already in the literature e.g. in [12, 5, 37].
We note that like other recent results on roots of sets of multivariate polynomials, e.g. [17], also the present paper is related to work in coding theory and cryptography. For this particular paper the inspiration came from work on ramp secret sharing schemes [24]. Elaborating further on this relationship we observe, that Theorem 17 should have interesting implications for one-point algebraic geometric codes [43, Sec. II] and other code constructions where one can obtain information on the functions $\bar{\sigma}$ as well as $\bar{\mu}$.